# Orthogonal Low-rank Adaptation in Lie Groups for Continual Learning of Large Language Models

### v1

@@ 5.2 Order of Taylor Expansion
In practice, to reduce computational overhead, we can approximate the exponential map using its Taylor expansion. While the resulting perturbation is not perfectly structure-preserving, it offers a close approximation. By expanding to a higher order, we can achieve a better fit to the ideal structure-preserving perturbation, analogous to approximating a continuous curve with a polynomial function. Consequently, we posit that the order of the Taylor expansion influences the final structure-preserving effect—higher orders yield approximations that more closely resemble the ideal infinitesimal perturbation that perfectly preserves structure. The mathematical expression for the $n$-th order Taylor expansion of the exponential map is as follows:
@@
where $I$ denotes the matrix of all ones, and $\Delta W^{\odot 2}$ represents the Hadamard (element-wise) product $\Delta W\odot\Delta W$.
To investigate this, we conducted an ablation study: on the Standard CL Benchmark, using the T5 base model, we trained models utilizing first-order, second-order, and third-order Taylor approximations, respectively. The final results, reported as the average over three independent runs, are presented in Table 3.
@@ 5.3 Effectiveness of the Multiplicative Update
To validate the necessity of the multiplicative update from the Lie group perspective, we conducted a critical ablation study. Under the condition of keeping the full subspace orthogonality constraint loss function unchanged, we replaced the proposed multiplicative parameter update ($W\odot\exp(\Delta W)$) with the standard additive update ($W+\Delta W$). This ablated model is denoted as "No LieGroup Mult". The experimental results are presented in Table 4.
A significant and consistent performance degradation (an average drop of approximately $2.3\%$) is observed across all task sequences of the Standard CL Benchmark when the multiplicative update mechanism is removed. This result strongly demonstrates that the exponential map-based multiplicative update adopted in this work is crucial for achieving optimal performance. It is not merely a replaceable implementation detail but a core component of the OLieRA method, working synergistically with the orthogonality constraint.
We hypothesize that the multiplicative update, due to its scaling nature, better preserves the intrinsic structure of the pre-trained weights. This provides a more stable and accommodating parameter space foundation for the subsequent orthogonally constrained updates for new tasks. Without this geometric structure-preserving mechanism, the effectiveness of the subspace orthogonality constraint is substantially diminished.
@@ 5.4 Unveiling the Internal Mechanism via Fisher Information
In continual learning, the Fisher Information Matrix is commonly used to quantify the importance of parameters for previously learned tasks. Given model parameters $\theta$, the $i$-th diagonal element of the Fisher Information Matrix is typically approximated empirically as:
$F_{i}\approx\mathbb{E}_{x\sim\mathcal{D}}\left[\left(\frac{\partial}{\partial\theta_{i}}\log p(y|x;\theta)\right)^{2}\right],$ (15)
where $\mathcal{D}$ is the data distribution of the old task and $\theta_{i}$ denotes the $i$-th parameter of the model. A larger $F_{i}$ indicates that the corresponding parameter direction is more sensitive to the performance of the old tasks, and thus should be more strongly protected during subsequent fine-tuning.
In LoRA-based continual learning methods, the learning process for each task introduces a parameter increment $\Delta\theta$. To quantify the degree of conflict between the updates for new and old tasks, we can define a Fisher-weighted energy metric for $\Delta\theta$ under the diagonal Fisher approximation Kirkpatrick et al. (2017):
$E=\sum_{i}F_{i}(\Delta\theta_{i})^{2}.$ (16)
This metric describes the "energy" of the new task’s update within the sensitive directions of the old tasks. Intuitively, a larger $E$ indicates that the parameter update for the new task causes stronger perturbations along directions important to the old tasks, posing a higher risk of catastrophic forgetting. Conversely, a smaller $E$ suggests that the new task’s update is more concentrated in directions with lower Fisher information, thereby having a relatively limited impact on the old tasks.
@@
We computed the Fisher-weighted energy of $\Delta\theta$ between the last and the penultimate tasks. The results are presented in Table 5. As shown, the Fisher Energy value of N-LoRA is close to zero, with a magnitude significantly smaller than that of O-LoRA, which is consistent with the observation in Yang et al. (2025) that N-LoRA achieves stronger orthogonality. Interestingly, in our experiments we found that OLieRA exhibits the opposite trend: its Fisher-weighted $\Delta\theta$ energy is higher than that of the other two baseline methods. This result suggests that, by employing exponential mapping and enforcing orthogonality in the Lie group parameter space, OLieRA does not merely avoid updates along the sensitive directions of previous tasks. Instead, it allows updates even in directions with large Fisher values. We argue that this strategy enhances parameter utilization, enabling new tasks to better leverage the important representations already learned. As a result, OLieRA achieves stronger expressiveness and superior overall performance, striking a balance between mitigating task interference and promoting knowledge sharing.
@@ A.2 Datasets
Table 6 provides detailed descriptions of the 15 datasets and their evaluation metrics used in our continual learning (CL) experiments. Overall, the datasets were drawn from three main benchmark suites: the CL benchmark Zhang et al. (2015), GLUE Wang et al. (2018), and SuperGLUE Wang et al. (2019). In addition, following Razdaibiedina et al. (2023), we supplemented these benchmarks with the IMDB movie reviews dataset.

### v2

@@ 5.2 Order of Taylor Expansion
To reduce computational overhead, we approximate the exponential map via a truncated Taylor expansion. Although this approximation is not strictly structure-preserving, higher-order expansions more closely match the ideal update, analogous to approximating a smooth curve with higher-degree polynomials. Hence, the expansion order directly controls the degree of structure preservation.
The $n$-th order Taylor expansion of the exponential map is
@@
where $I$ is the all-ones matrix and $\Delta W^{\odot 2}=\Delta W\odot\Delta W$ denotes the Hadamard product.
We validate this effect through an ablation study on the Standard CL Benchmark with T5-base, comparing first-, second-, and third-order approximations. Results averaged over three runs are reported in Table 3.
@@ 5.3 Effectiveness of the Multiplicative Update
To assess the necessity of the Lie group–motivated multiplicative update, we perform an ablation study in which the proposed update $W\odot\exp(\Delta W)$ is replaced with a standard additive update $W+\Delta W$, while keeping the subspace orthogonality constraint unchanged. We refer to this variant as No LieGroup Mult. Results are reported in Table 4.
Removing the multiplicative update leads to a consistent and substantial performance drop across all task sequences on the Standard CL Benchmark, with an average degradation of approximately $2.3\%$. This confirms that the exponential map–based multiplicative update is a core component of OLieRA rather than a replaceable implementation detail, and that it works synergistically with the orthogonality constraint.
We attribute this effect to the scaling nature of the multiplicative update, which better preserves the intrinsic structure of pre-trained weights and provides a more stable parameter space for subsequent orthogonally constrained adaptations. Without this structure-preserving mechanism, the benefit of the subspace orthogonality constraint is significantly reduced.
@@ 5.4 Unveiling the Internal Mechanism via Fisher Information
In continual learning, the Fisher Information Matrix (FIM) is widely used to measure the importance of parameters for previously learned tasks. For model parameters $\theta$, the $i$-th diagonal element is empirically approximated as
$F_{i}\approx\mathbb{E}_{x\sim\mathcal{D}}\left[\left(\partial_{\theta_{i}}\log p(y\mid x;\theta)\right)^{2}\right],$ (18)
where larger $F_{i}$ indicates greater sensitivity of old-task performance to $\theta_{i}$, and thus stronger protection during fine-tuning.
In LoRA-based continual learning, each task induces a parameter update $\Delta\theta$. Under the diagonal FIM approximation Kirkpatrick et al. (2017), the conflict between new and old tasks can be quantified by the Fisher-weighted energy
$E=\sum_{i}F_{i}(\Delta\theta_{i})^{2},$ (19)
which measures how strongly the new update perturbs directions important to previous tasks.
@@
We compute this energy between the last two tasks (Table 5). Consistent with prior work Yang et al. (2025), N-LoRA yields near-zero energy, indicating strong orthogonality, while O-LoRA produces larger values. In contrast, OLieRA exhibits higher Fisher-weighted energy than both baselines, suggesting that its Lie group–based updates do not simply avoid sensitive directions but enable controlled updates along them. We argue that this improves parameter utilization and knowledge sharing, resulting in greater expressiveness and superior overall performance despite stronger interaction with high-Fisher directions.
@@ 6 Conclusion
7 Ethics and AI Disclosure
We used ChatGPT to assist with text drafting. All scientific content and experiments were independently done by the authors.
@@ A.2 Datasets
Table 6 provides detailed descriptions of the 15 datasets and their evaluation metrics used in our continual learning (CL) experiments. Overall, the datasets were drawn from three main benchmark suites: the CL benchmark Zhang et al. (2015), GLUE Wang et al. (2018), and SuperGLUE Wang et al. (2019). In addition, following Razdaibiedina et al. (2023), we supplemented these benchmarks with the IMDB movie reviews dataset. Data is secure and has been anonymized.
@@ A.4 Task Instructions
A.5 Limitations
Despite its empirical effectiveness, OLieRA has several limitations that warrant further investigation.
First, the proposed formulation models model parameters as elements of an Abelian Lie group defined by the Hadamard product. While this choice enables efficient and stable multiplicative updates, it remains unclear to what extent faithfully reflects structures that may underlie large language model parameters.
Second, compared with strong baselines such as O-LoRA and N-LoRA, the performance improvement brought by our proposed method is relatively modest in long task sequences, which may imply that more sophisticated approaches are required to maintain the parameter structure for long task sequences.
Finally, the method introduces additional hyperparameters (e.g., the strength of orthogonality regularization), whose tuning may be nontrivial and task-order dependent.
Addressing these limitations, particularly through richer geometric parameterizations and more principled approximation schemes, is an important direction for future work.
Potential risks of our continual learning approach include performance bias across tasks or populations, privacy concerns when models are trained on sensitive data, and possible misuse in high-stakes decision settings if the model is applied outside controlled experiments.